Session: Jun 2012 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Tue Jun 05 06:29:42 2012) Failure: ANT
Observations could not be carried due to serious problems with telescope control system which arose before this experiment started. Although this could be premature diagnosis, probably the problems are connected with recent upgrade of the angle encoder in azimuth to a higher resolution one.

Medicina (M. Roma at Tue Jun 05 07:07:16 2012) Success:
All scan recorded. sky clear.

Yebes (Daniel Sastre at Tue Jun 05 07:08:28 2012) Success:

Urumqi (Peng Li at Tue Jun 05 07:53:44 2012) Success:

Westerbork (Geert Kuper at Tue Jun 05 07:57:04 2012) Success:
Source byond HA limits until 03:08 UT

Hartebeesthoek (Jonathan Quick at Tue Jun 05 08:00:12 2012) Success:
No known problems

Noto (P. Cassaro at Tue Jun 05 15:44:41 2012) Success:

Badary (Holkina at Thu Jun 07 12:21:01 2012) Success:

Zelenchukskaya (Ptitcin A. and Kravchenko G. at Thu Jun 07 12:21:22 2012) Success:

Svetloe (Shumilova at Thu Jun 07 12:22:29 2012) Minor failures:RECE
warm receiver

Onsala (Michael Lindqvist at Thu Jun 07 13:58:21 2012) Minor failures:
A few Mark5 warnings (see log), otherwise no known problems.

Noto (P. Cassaro at Mon Jun 11 10:04:07 2012) Success:

Effelsberg (Uwe Bach at Wed Jun 13 09:07:38 2012) Success: